OpenStudy (anonymous):

HELP ASAP MEDAL! A rectangle has a width length of 5sqrt8 inches and a width of 2sqrt8 inches. What is the perimeter? I got 56 inches but I dont think that is right.

OpenStudy (anonymous):

do you mean \[2^{8}\] and \[5^{8}\] ?

OpenStudy (anonymous):

No, 2\[2\sqrt{8} \] and the same for the other one

OpenStudy (anonymous):

\[5\sqrt{8}\]

OpenStudy (anonymous):

oh okay give me a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[2\sqrt{8}\] = 5.65 \[5\sqrt{8}\] = 14.14 if rounding = 40 if not = 38 hope that helped

OpenStudy (anonymous):

ok so would it be \[14\sqrt{8}\]

OpenStudy (anonymous):

there arent any options for that if it would be 38 or 40..

OpenStudy (anonymous):

Pretty sure its the 14sqrt8 one

OpenStudy (anonymous):

Yes it is! Sorry i'm still in school Dx